# Randomized Black-Box PIT for Small Depth +-Regular Non-commutative Circuits

### v1

@@ 5.2 Randomized Identity Test for Small Depth +-Regular Circuits
We use Thoerem 6 to convert the n.c. polynomial $f$ into an $s^{(d^{+}-1)}$-ordered power-sum polynomial $f_{ops}$, while preserving its non-zeroness. As discussed above, there is a substitution automaton of size bounded by $s^{O(d^{2})}$, which results in substitution matrices of dimension $s^{O(d^{2})}$. By Claim 4.0.1, $f_{ops}$ can be treated as a commutative polynomial while preserving its non-zeroness. Using the DeMillo-Lipton-Schwartz-Zippel lemma, we can have a randomized PIT for depth $d$ $+$-regular circuit of size $s$ using matrices of dimension at most $s^{O(d^{2})}$. This completes the proof of the theorem. ∎
@@ Acknowledgements
    Replace the variable $x_{i}$ variable that appears in $[D_{1}]\setminus I$ with a new commuting variable $z_{i}$.
@@
Here, we will present the proof of the product sparsification lemma. This proof follows the proof of Lemma 6.2 in [AJMR19], but crucially uses Claim 4.0.1.
@@
  • Induction hypothesis: Assume the lemma holds for sums of products of $s$-ordered power-sum polynomials with at most $s-1$ terms.
@@
    Now, we prove the forward direction. Let $P_{i}=\prod_{j\in[D_{2}]}\hat{Q}_{ij}$ for all $i\in[s]$. Suppose $\sum_{i\in[s]}\beta_{i}P_{i}\neq 0$. We want to show there exists a subset $J\subseteq[D_{2}]$ of size at most $s-1$ such that $\sum_{i\in[s]}\beta_{i}P_{i,J}\neq 0$, where $P_{i,J}=\left(\prod_{j\in\overline{J}}\hat{Q}_{ij}\right)\left(\prod_{j\in J}\hat{Q}_{ij}\right)$.
@@
    $\sum^{s}_{i=1}\beta_{i}P_{i}=P\times\left(\sum^{t}_{i=1}\beta_{i}Q_{i,j_{0}}P^{\prime}_{i}\right)+P\times\left(\sum^{s}_{i=t+1}\beta_{i}Q_{i,j_{0}}P^{\prime}_{i}\right).$
@@
    $P\times\left(\sum^{s}_{i=t+1}\beta_{i}Q_{i,j_{0}}P^{\prime}_{i}\right)=P\times\left(\sum^{s}_{i=t+1}\beta_{i}\left(\sum_{\ell\in[t]}\alpha^{(i)}_{\ell}\hat{Q}_{\ell,j_{0}}\right)P^{\prime}_{i}\right).$
    $\sum^{s}_{i=1}\beta_{i}P_{i}$ $=$ $P\times\left(\sum^{t}_{i=1}\beta_{i}Q_{i,j_{0}}P^{\prime}_{i}\right)+P\times\left(\sum^{s}_{i=t+1}\beta_{i}\left(\sum_{\ell\in[t]}\alpha^{(i)}_{\ell}\hat{Q}_{\ell,j_{0}}\right)P^{\prime}_{i}\right)$
    $=$ $P\times\sum^{t}_{k=1}Q_{k,j_{0}}P^{{}^{\prime\prime}}_{k}$
@@
    Since we treat all polynomials $Q_{k,j_{0}},k\in[t]$ as n.c. and all $Q_{ij}$ are homgeneous polynomials of the same degree $D_{1}$, with each $Q_{k,j_{0}}$ being $s$-ordered power-sum polynomials, we can view them as linear forms as follows. This is only for analysis. Let $Y=\{y_{\overline{\ell}}\mid\ell=(\ell_{1},\ldots,\ell_{s})\text{ and }\sum_{i}\ell_{i}=D_{1}\}$ be the set of new non-commuting variables. Each monomial $m_{a}$ with coefficient $\beta_{m_{a}}$ in $Q_{k,j_{0}}$ can be expressed as $m_{a}=\xi_{1}^{a_{1}}\cdot\xi_{2}^{a_{2}}\cdots\xi_{s}^{a_{s}}$ and can represented by $\left(\beta_{m_{a}}\cdot z_{\overline{a}}\right)$. Thus, we can envision $Q_{k,j_{0}}$ polynomial as a linear form over $Y$ variables, denoted by $L_{k,j_{0}}$. Note that there is a bijection between monomials of $Q_{k,j_{0}}$ and $L_{k,j_{0}}$ (which consists of only variables from $Y$).
    Since all polynomials $Q_{k,j_{0}}$ for $k\in[t]$ are linearly independent, it follows that the linear forms $L_{k,j_{0}},k\in[t]$ are also linearly independent.
@@ C.1 Substitution Automaton for Commutative Transformation
Recall that $\tilde{f}=\hat{f}+F\in\mathbb{F}\mathopen{\langle}\xi\mathclose{\rangle}$ is the polynomial obtained after Step (1). It is important to note that any monomial in $\tilde{f}$, including all monomials in $A^{\tilde{m}}_{{\hat{f}}}$ and $B_{F}^{\tilde{m}}$, is formed by concatenating $\xi$-patterns (see Defintion 4.0.2).
Consider a monomial $m_{j}=m_{j,1}m_{j,2}\cdots m_{j,N_{j}}\in B_{F}^{\tilde{m}}$ , where $m_{j,1},m_{j,2},\ldots,m_{j,N_{j}}$ are $\xi$-patterns. While the degree of $m_{j}$ matches the overall degree $D=D_{1}\times D_{2}$ of the polynomial, we cannot assert whether $N_{j}$ is equal to $D_{2}$ or not (See Case 2 in 4.1). However, we can affirm that there is at least one $\xi$-pattern in $m_{j}$ for which the sum of the exponents does not equal $D_{1}$.
For $B_{F}^{\tilde{m}}$, there are two possibilities:
@@
$\tilde{f}(M_{\xi_{1}},M_{\xi_{2}},\ldots,M_{\xi_{s}})=\hat{f}(M_{\xi_{1}},M_{\xi_{2}},\ldots,M_{\xi_{s}})+F(M_{\xi_{1}},M_{\xi_{2}},\ldots,M_{\xi_{s}}).$
@@
  • Case 1: There exists a monomial $m_{l}\in B_{F}^{\tilde{m}}$ such that $N_{l}\neq D_{2}$. Let $\lambda\equiv D_{2}\mod p$.
@@
    Let’s fix such a monomial $m_{l}\in B_{F}^{\tilde{m}}$ represented as $m_{\ell}=m_{\ell,1}m_{\ell,2}\cdots m_{\ell,N_{\ell}}$. As previously noted, an automaton can identify the boundaries between any two sub-monomials $m_{\ell,i}$ and $m_{\ell,i+1}$ (for $i<N_{\ell}$) in the monomial $m_{\ell}$. Therefore, for any natural number $p$ we can compute $N_{\ell}\mod p$ using an automaton.
@@
    Recall that for each monomial in $A^{\tilde{m}}_{{\hat{f}}}$, the number of sub-monomials (i.e., the number of $\xi$-patterns) is exactly $D_{2}$. We can compute the number of $\xi$-patterns modulo $p$ as boundaries between any two sub-monomials $m_{\ell,i}$ and $m_{\ell,i+1}$ can be identified by the automaton.
@@
    Since $N_{j},D_{2}\leq D$, by Proposition D.0.1, there exists an automaton $\mathcal{M}$ with a number of states bounded by $O(\log D)$. If we consider the output state of $\mathcal{M}$ when processing $m_{l}$, it will differ from the output state of all monomials in $A^{\tilde{m}}_{{\hat{f}}}$. The automaton $\mathcal{M}$ does not change the monomial; it simply maps the n.c. variable $\xi_{i}$ to itself while computing the number of $\xi$-patterns modulo $p$. Notably, all monomials in $A^{\tilde{m}}_{{\hat{f}}}$ reach the same state in the automaton, which we denote as $\lambda$ (with $\lambda=O(\log D)$).
@@
    By Proposition D.0.1, at least one monomial $m_{l}\in B_{F}^{\tilde{m}}$ will not reach the state $\lambda$. If we consider the output of the automaton $\mathcal{M}$ in Figure 4 on $\tilde{f}$ as $(q_{0},q_{\lambda})$-th entry of the resulting matrix, then the n.c. polynomial at this entry is given by
    $f_{\lambda}=\hat{f}+\tilde{F}$
@@
    where the coefficient of the monomial $m_{l}$ in the polynomial $\tilde{F}$ is 0. Importantly, since $m_{l}\in B_{F}^{\tilde{m}}$ the coefficient of $m_{l}$ in $F$ is non-zero.
@@
    It is crucial to note that $f_{\lambda}\neq 0$ because $\hat{f}\neq 0$ (by Lemma 4.1.1), and the non-zero monomials of $\tilde{F}$ are a subset of the non-zero monomials of $F$, with no common non-zero monomials between $\hat{f}$ and $F$ (see Claims 4.1.1 and 4.1.2).
@@
    Additionally, while $\tilde{f}\neq f_{\lambda}$, all monomials of $A^{\tilde{m}}_{{\hat{f}}}$ are included in $f_{\lambda}$ with the same coefficients as in $\tilde{f}$. However, at least one monomial $m_{l}\in B_{F}^{\tilde{m}}$ is absent from $\tilde{F}$, thus from $f_{\lambda}$. Since only monomials in $A^{\tilde{m}}_{{\hat{f}}}\sqcup B_{F}^{\tilde{m}}$ are transformed into the commutative monomial $\tilde{m}$ during product sparsification and commutative transformation (Steps 2 and 3), performing these steps on $f_{\lambda}$ will yield a non-zero commutative polynomial, specifically ensuring that the coefficient of $\tilde{m}$ is non-zero in the transformed commutative polynomial.
  • Case 2: For each monomial $m_{\ell}\in B_{m}^{F}$, $N_{\ell}=D_{2}$. Let $\lambda\equiv D_{1}\mod p$.
    Let’s fix a monomial $m_{\ell}\in B_{m}^{F}$ expressed as $m_{\ell}=m_{\ell,1}m_{\ell,2}\cdots m_{\ell,D_{2}}$. For any monomial $m_{\ell}$ and $r\in[D_{2}]$, we define
@@
    We know that there exists an $\xi$-pattern $m_{\ell,r}$ in the monomial $m_{\ell}$, such that $\mathcal{K}_{m_{\ell,r}}\neq D_{1}$ by Claim 4.1.2. For all $m_{t}\in A^{\tilde{m}}_{{\hat{f}}}$, it follows that $\mathcal{K}_{m_{t,r}}=D_{1}$ for each $r\in[D_{2}]$ (as $N_{t}=D_{2}$) by Claim 4.1.1.
@@
    Given that $\mathcal{K}_{m_{\ell,r}},D_{1}\leq D$, by Proposition D.0.1, there exists an automaton $\mathcal{M}$ that computes $\left(\mathcal{K}_{m_{\ell,r}}\mod p\right)$. The automaton we construct has the number of states bounded by $O(s^{2})$, that is $O(\log^{2}D)$ as $D$ is exponential in $s$ for us. The output state of $\mathcal{M}$ when processing $m_{l}$ will differ from the output state of all monomials in $A^{\tilde{m}}_{{\hat{f}}}$.
@@
    For any monomial $m_{t}\in A^{\tilde{m}}_{{\hat{f}}}$, for every guess $r\in[D_{2}]$, we have $\lambda\equiv\mathcal{K}_{m_{t,r}}\mod p$, where $\mathcal{K}_{m_{t,r}}=D_{1}$. Thus, when considering the output of the automaton $\mathcal{M}$ on the monomial $m_{\ell}$ as the sum of two entries
@@
    Conversely, for the above fixed monomial $m_{\ell}\in B_{F}^{\tilde{m}}$, there exists a guess $r\in[D_{2}]$ such that $\mathcal{K}_{m_{\ell,r}}\neq D_{1}$ (by Claim 4.1.2). Therefore, this computation path will not reach either of the final states $q_{f_{1}}$ or $q_{f_{2}}$. Thus, the coefficient $\alpha_{m_{\ell}}$ of the monomial $m_{\ell}$ is scaled by at most a factor of $(D_{2}-1)$ in $M[q_{0},q_{f_{1}}]+M[q_{0},q_{f_{2}}]$.
@@
    $f_{\lambda}=D_{2}\cdot\hat{f}+\tilde{F},$

### v2

@@ 5.2 Randomized Identity Test for Small Depth +-Regular Circuits
We use Theorem 6 to convert the n.c. polynomial $f$ into an $s^{(d^{+}-1)}$-ordered power-sum polynomial $f_{ops}$, while preserving its non-zeroness. As discussed above, there is a substitution automaton of size bounded by $s^{O(d^{2})}$, which results in substitution matrices of dimension $s^{O(d^{2})}$. By Claim 4.0.1, $f_{ops}$ can be treated as a commutative polynomial while preserving its non-zeroness. Using the DeMillo-Lipton-Schwartz-Zippel lemma, we can have a randomized PIT for depth $d$ $+$-regular circuit of size $s$ using matrices of dimension at most $s^{O(d^{2})}$. This completes the proof of the theorem. ∎
@@ Acknowledgement
    Replace the variable $x_{i}$ that appears in $[D_{1}]\setminus I$ with a new commuting variable $z_{i}$.
@@
A.1
Proof of Lemma 5.1.1
Lemma A.1.1.
Let
$f=\sum_{i\in[s]}\prod_{j\in[D_{2}]}Q_{ij}$
be a non-commutative polynomial over $X=\{x_{1},\cdots,x_{n}\}$, computed by a depth $d$ $+$-regular circuit of size $s$, where each $Q_{ij}$ is computed by a depth $(d-2)$ $+$-regular circuit of size at most $s$. Let $D_{1}\geq s-1$ denote the degree of the polynomial $Q_{ij}$, for $i\in[s],j\in[D_{2}]$.
Suppose that any non-commutative polynomial $P$, computed by a depth $(d-2)$ $+$-regular circuit, can be converted into a non-commutative polynomial $\hat{P}$ (which is an ordered power- sum polynomial) while preserving non-zeroness. Then, each $Q_{ij}$ polynomial in $f$ can be converted to $\hat{Q}_{ij}$ to create the n.c. polynomial $\hat{f}=\sum_{i\in[s]}\prod_{j\in[D_{2}]}\hat{Q}_{ij}$. Then, if $f\neq 0$ then $\hat{f}\neq 0$ maintaining the non-zeroness of the entire polynomial $\hat{f}$. That is, $f=0$ if and only if $\hat{f}=0$.
Proof.
The proof is by induction on $D_{2}$. Assume $f$ is non-zero.
  • Base Case: $D_{2}=1$. That is, the number of terms in each product is 1, and we have
    $f=\sum_{i\in[s]}Q_{i1}.$
    For $D_{2}=1$, each $Q_{i1}$ for $i\in[s]$ can be computed by a depth-$(d-2)$ $+$-regular circuit. Consequently, the polynomial $f$ can be computed by a depth-$(d-2)$ $+$-regular circuit of size at most $s$. This is possible because the top gate of each circuit computing $Q_{i1}$ is a sum gate (denoted as a $\Sigma$-gate), and the sum gates across all circuits can be combined into a single $\Sigma$-gate without increasing the depth of the circuit. Therefore, the overall circuit computing $\sum_{i\in[s]}Q_{i1}$ retains a depth of $d-2$.
    Using one of the hypotheses of the lemma: if any polynomial $P$, computed by a depth $(d-2)$ $+$-regular circuit, can be converted into a $\hat{P}$ (which is an ordered power-sum polynomial) while preserving non-zeroness, we conclude that the polynomial $f$ can be transformed into another polynomial $\hat{f}$ preserving non-zeroness. This completes the proof of the base case.
  • Inductive Step:
    Consider the polynomial $f$ expressed as:
    $f=\sum_{i\in[s]}Q_{i1}P_{i},$
    where $D=D_{1}\times D_{2}$ is the degree of $f$ and $\left([m]P_{i}\right)\in\mathbb{F}$. Since $f\not\equiv 0$, there exists a monomial $m\in X^{D-D_{1}}$ such that the right derivative $f^{(m)}$ does not vanish. We define $f^{(m)}$ as:
    $f^{(m)}=\sum_{i\in[s]}Q_{i1}\times\left([m]P_{i}\right),$
    where $\left([m]P_{i}\right)\in\mathbb{F}$.
    The polynomial $f^{(m)}\not\equiv 0$ can be computed by a depth $(d-2)$ +-regular circuit of size at most $s$. This is because as each $Q_{i1}$ can be computed by a depth $(d-2)$ +-circuit, then the sum $\sum_{i\in[s]}Q_{i1}\times\left([m]P_{i}\right)$ can also be computed by a depth-($d-2$) $+$-regular circuit. This is because the top gate of each circuit computing $Q_{i1}$ is a sum gate (denoted as a $\Sigma$-gate), and the sum gates across all circuits can be combined into a single $\Sigma$-gate without increasing the depth of the circuit. Therefore, the depth of the overall circuit computing $\sum_{i\in[s]}Q_{i1}\times\left([m]P_{i}\right)$ remains $d-2$.
    This reduces to the base case. Therefore, $f^{(m)}$ can be transformed into
    $\hat{f}^{(m)}=\sum_{i\in[s]}\hat{Q}_{i1}\times\left([m]P_{i}\right)$
    while preserving non-zeroness. Since $\hat{f}^{(m)}$ is non-zero, the polynomial
    $f^{{\dagger}}=\sum_{i\in[s]}\hat{Q}_{i1}P_{i},$
    Note that the polynomial $\hat{Q}_{i1}$ is an ordered power-sum non-commutative polynomial over some new set of non-commutative variables. Now, expand the polynomial $\hat{Q}_{i1}$ as a sum of monomials in $f^{{\dagger}}$:
    Since $f^{{\dagger}}\not\equiv 0$, there exists a monomial $m$ such that the left derivative of $f^{{\dagger}}$ with respect to $m$ does not vanish. Let $f^{{\dagger}}_{(m)}$ be this polynomial. Again, note that this derivative is for analysis only, and we will not compute it. Let $\beta_{i}=[m]\hat{Q}_{i1}$. Note that $\beta_{i}\in\mathbb{F}$.
    $f^{{\dagger}}_{(m)}=\sum_{i\in[s]}\beta_{i}\times P_{i}=\sum_{i\in[s]}\beta_{i}\times\prod_{j=2}^{D_{2}}Q_{ij}.$
    Clearly, $f^{{\dagger}}_{(m)}\not\equiv 0$ and the number of product terms in the polynomial $P_{i}$ is exactly $(D_{2}-1)$.
    By induction hypothesis, each $Q_{ij}$ in $f^{{\dagger}}_{(m)}$ can be transformed into $\hat{Q}_{ij}$ polynomial such that the resulting polynomial is non-zero. Since $f^{{\dagger}}_{(m)}\not\equiv 0$, the following polynomial is also non-zero:
    $\hat{f}=\sum_{i\in[s]}\prod_{j\in[D_{2}]}\hat{Q}_{ij}.$
@@
We recall the following proposition from [AJMR19] (see Proposition 3.1).
Proposition B.0.1 ([AJMR19]).
Let $A:\mathbb{F}^{n}\to\mathbb{F}^{n}$ be an invertible linear transformation, and let $f(x_{1},x_{2},\dots,x_{n})\in\mathbb{F}\langle X\rangle$ be any homogeneous polynomial of degree $d$. Let $A_{j}(f)$ be the polynomial obtained by applying the transform $A$ to the $j$-th position of monomials in $f$ for $j\in[d]$. Then, $f\neq 0$ if and only if $A_{j}(f)\neq 0$.
Here, we will present the proof of the product sparsification lemma. This proof follows the proof of Lemma 6.2 in [AJMR19], but crucially uses Claim 4.0.1. Unlike [AJMR19], the key distinction in our setting is that the $\hat{Q}_{ij}$ polynomial can be non-homogeneous.
@@
  • Induction hypothesis: Assume the lemma holds for sums of products of $s$-ordered power-sum polynomials with at most $s-1$ terms. Recall
    $\hat{f}_{1}=\sum_{i\in[s]}\prod_{j\in[D_{2}]}\hat{Q}_{ij},$
    where each $\hat{Q}_{ij}$ is an $s$-ordered power-sum polynomial of degree $D_{1}$ over $\xi=\{\xi_{1},\xi_{2},\ldots,\xi_{s}\}$.
@@
    Now, we prove the forward direction.
    It is important to note that the $\hat{Q}_{ij}$ polynomials can, in general, be non-homogeneous (but in depth-5 case, each $\hat{Q}_{ij}$ is a homogeneous polynomial). However, each product within the summand involves the same number of $\hat{Q}_{ij}$ polynomials. Specifically, each summand is a product of exactly $D_{2}$ $\hat{Q}_{ij}$ polynomials. Therefore, we will treat the $\hat{Q}_{ij}$ polynomials as a whole, rather than expanding them as a sum of monomials or multiplying them with other $\hat{Q}_{ij}$ polynomials in the product. This approach is valid because the boundaries of the $\hat{Q}_{ij}$ polynomials can be efficiently identified, allowing us to select the entire $\hat{Q}_{ij}$ polynomial as either non-commutative or commutative.
    Let $P_{i}=\prod_{j\in[D_{2}]}\hat{Q}_{ij}$ for all $i\in[s]$. Suppose $\sum_{i\in[s]}\beta_{i}P_{i}\neq 0$. We want to show there exists a subset $J\subseteq[D_{2}]$ of size at most $s-1$ such that $\sum_{i\in[s]}\beta_{i}P_{i,J}\neq 0$, where $P_{i,J}=\left(\prod_{j\in\overline{J}}\hat{Q}_{ij}\right)\left(\prod_{j\in J}\hat{Q}_{ij}\right)$.
@@
    $\sum^{s}_{i=1}\beta_{i}P_{i}=P\times\left(\sum^{t}_{i=1}\beta_{i}\hat{Q}_{i,j_{0}}P^{\prime}_{i}\right)+P\times\left(\sum^{s}_{i=t+1}\beta_{i}\hat{Q}_{i,j_{0}}P^{\prime}_{i}\right).$
@@
    $P\times\left(\sum^{s}_{i=t+1}\beta_{i}\hat{Q}_{i,j_{0}}P^{\prime}_{i}\right)=P\times\left(\sum^{s}_{i=t+1}\beta_{i}\left(\sum_{\ell\in[t]}\alpha^{(i)}_{\ell}\hat{Q}_{\ell,j_{0}}\right)P^{\prime}_{i}\right).$
    $\sum^{s}_{i=1}\beta_{i}P_{i}$ $=$ $P\times\left(\sum^{t}_{i=1}\beta_{i}\hat{Q}_{i,j_{0}}P^{\prime}_{i}\right)+P\times\left(\sum^{s}_{i=t+1}\beta_{i}\left(\sum_{\ell\in[t]}\alpha^{(i)}_{\ell}\hat{Q}_{\ell,j_{0}}\right)P^{\prime}_{i}\right)$
    $=$ $P\times\sum^{t}_{k=1}\hat{Q}_{k,j_{0}}P^{{}^{\prime\prime}}_{k}$
@@
    Since we treat all polynomials $\hat{Q}_{k,j_{0}},k\in[t]$ as n.c with each $\hat{Q}_{k,j_{0}}$ being $s$-ordered power-sum polynomials, we can view them as linear forms as follows. This is only for analysis. Let
    $Y=\{y_{\overline{\ell}}\mid\ell=(\ell_{1},\ldots,\ell_{s})\text{ and }\sum_{i}\ell_{i}\leq D\}$
    be the set of new non-commuting variables, where $D$ is the degree of the given polynomial $\hat{f}_{1}$. Each monomial $m_{a}$ with coefficient $\beta_{m_{a}}$ in $\hat{Q}_{k,j_{0}}$ can be expressed as $m_{a}=\xi_{1}^{a_{1}}\cdot\xi_{2}^{a_{2}}\cdots\xi_{s}^{a_{s}}$ and can represented by $\left(\beta_{m_{a}}\cdot y_{\overline{a}}\right)$. Thus, we can envision $\hat{Q}_{k,j_{0}}$ polynomial as a linear form over $Y$ variables, denoted by $L_{k,j_{0}}$. Note that there is a bijection between monomials of $\hat{Q}_{k,j_{0}}$ and $L_{k,j_{0}}$ (which consists of only variables from $Y$).
    Since all polynomials $\hat{Q}_{k,j_{0}}$ for $k\in[t]$ are linearly independent, it follows that the linear forms $L_{k,j_{0}},k\in[t]$ are also linearly independent.
@@ C.1 Substitution Automaton for Commutative Transformation
Recall that $\tilde{f}=\hat{f}_{1}+F_{1}\in\mathbb{F}\mathopen{\langle}\xi\mathclose{\rangle}$ is the polynomial obtained after Step (1). It is important to note that any monomial in $\tilde{f}$, including all monomials in $A^{\tilde{m}}_{{\hat{f}_{1}}}$ and $B_{F_{1}}^{\tilde{m}}$, is formed by concatenating $\xi$-patterns (see Defintion 4.0.2).
Consider a monomial $m_{j}=m_{j,1}m_{j,2}\cdots m_{j,N_{j}}\in B_{F_{1}}^{\tilde{m}}$ , where $m_{j,1},m_{j,2},\ldots,m_{j,N_{j}}$ are $\xi$-patterns. While the degree of $m_{j}$ matches the overall degree $D=D_{1}\times D_{2}$ of the polynomial, we cannot assert whether $N_{j}$ is equal to $D_{2}$ or not (See Case 2 in 4.1). However, we can affirm that there is at least one $\xi$-pattern in $m_{j}$ for which the sum of the exponents does not equal $D_{1}$.
For $B_{F_{1}}^{\tilde{m}}$, there are two possibilities:
@@
$\tilde{f}(M_{\xi_{1}},M_{\xi_{2}},\ldots,M_{\xi_{s}})=\hat{f}_{1}(M_{\xi_{1}},M_{\xi_{2}},\ldots,M_{\xi_{s}})+F_{1}(M_{\xi_{1}},M_{\xi_{2}},\ldots,M_{\xi_{s}}).$
@@
  • Case 1: There exists a monomial $m_{l}\in B_{F_{1}}^{\tilde{m}}$ such that $N_{l}\neq D_{2}$. Let $\lambda\equiv D_{2}\mod p$.
@@
    Let’s fix such a monomial $m_{l}\in B_{F_{1}}^{\tilde{m}}$ represented as $m_{\ell}=m_{\ell,1}m_{\ell,2}\cdots m_{\ell,N_{\ell}}$. As previously noted, an automaton can identify the boundaries between any two sub-monomials $m_{\ell,i}$ and $m_{\ell,i+1}$ (for $i<N_{\ell}$) in the monomial $m_{\ell}$. Therefore, for any natural number $p$ we can compute $N_{\ell}\mod p$ using an automaton.
@@
    Recall that for each monomial in $A^{\tilde{m}}_{{\hat{f}_{1}}}$, the number of sub-monomials (i.e., the number of $\xi$-patterns) is exactly $D_{2}$. We can compute the number of $\xi$-patterns modulo $p$ as boundaries between any two sub-monomials $m_{\ell,i}$ and $m_{\ell,i+1}$ can be identified by the automaton.
@@
    Since $N_{j},D_{2}\leq D$, by Proposition D.0.1, there exists an automaton $\mathcal{M}$ with a number of states bounded by $O(\log D)$. If we consider the output state of $\mathcal{M}$ when processing $m_{l}$, it will differ from the output state of all monomials in $A^{\tilde{m}}_{{\hat{f}_{1}}}$. The automaton $\mathcal{M}$ does not change the monomial; it simply maps the n.c. variable $\xi_{i}$ to itself while computing the number of $\xi$-patterns modulo $p$. Notably, all monomials in $A^{\tilde{m}}_{{\hat{f}_{1}}}$ reach the same state in the automaton, which we denote as $\lambda$ (with $\lambda=O(\log D)$).
@@
    By Proposition D.0.1, at least one monomial $m_{l}\in B_{F_{1}}^{\tilde{m}}$ will not reach the state $\lambda$. If we consider the output of the automaton $\mathcal{M}$ in Figure 4 on $\tilde{f}$ as $(q_{0},q_{\lambda})$-th entry of the resulting matrix, then the n.c. polynomial at this entry is given by
    $f_{\lambda}=\hat{f}_{1}+\tilde{F}_{1}$
@@
    where the coefficient of the monomial $m_{l}$ in the polynomial $\tilde{F}_{1}$ is 0. Importantly, since $m_{l}\in B_{F_{1}}^{\tilde{m}}$ the coefficient of $m_{l}$ in $F$ is non-zero.
@@
    It is crucial to note that $f_{\lambda}\neq 0$ because $\hat{f}_{1}\neq 0$ (by Lemma 4.1.1), and the non-zero monomials of $\tilde{F}_{1}$ are a subset of the non-zero monomials of $F_{1}$, with no common non-zero monomials between $\hat{f}_{1}$ and $F_{1}$ (see Claims 4.1.1 and 4.1.2).
@@
    Additionally, while $\tilde{f}\neq f_{\lambda}$, all monomials of $A^{\tilde{m}}_{{\hat{f}_{1}}}$ are included in $f_{\lambda}$ with the same coefficients as in $\tilde{f}$. However, at least one monomial $m_{l}\in B_{F_{1}}^{\tilde{m}}$ is absent from $\tilde{F}_{1}$, thus from $f_{\lambda}$. Since only monomials in $A^{\tilde{m}}_{{\hat{f}_{1}}}\sqcup B_{F_{1}}^{\tilde{m}}$ are transformed into the commutative monomial $\tilde{m}$ during product sparsification and commutative transformation (Steps 2 and 3), performing these steps on $f_{\lambda}$ will yield a non-zero commutative polynomial, specifically ensuring that the coefficient of $\tilde{m}$ is non-zero in the transformed commutative polynomial.
  • Case 2: For each monomial $m_{\ell}\in B_{F_{1}}^{\tilde{m}}$, $N_{\ell}=D_{2}$. Let $\lambda\equiv D_{1}\mod p$.
    Let’s fix a monomial $m_{\ell}\in B_{F_{1}}^{\tilde{m}}$ expressed as $m_{\ell}=m_{\ell,1}m_{\ell,2}\cdots m_{\ell,D_{2}}$. For any monomial $m_{\ell}$ and $r\in[D_{2}]$, we define
@@
    We know that there exists an $\xi$-pattern $m_{\ell,r}$ in the monomial $m_{\ell}$, such that $\mathcal{K}_{m_{\ell,r}}\neq D_{1}$ by Claim 4.1.2. For all $m_{t}\in A^{\tilde{m}}_{{\hat{f}_{1}}}$, it follows that $\mathcal{K}_{m_{t,r}}=D_{1}$ for each $r\in[D_{2}]$ (as $N_{t}=D_{2}$) by Claim 4.1.1.
@@
    Given that $\mathcal{K}_{m_{\ell,r}},D_{1}\leq D$, by Proposition D.0.1, there exists an automaton $\mathcal{M}$ that computes $\left(\mathcal{K}_{m_{\ell,r}}\mod p\right)$. The automaton we construct has the number of states bounded by $O(s^{2})$, that is $O(\log^{2}D)$ as $D$ is exponential in $s$ for us. The output state of $\mathcal{M}$ when processing $m_{l}$ will differ from the output state of all monomials in $A^{\tilde{m}}_{{\hat{f}_{1}}}$.
@@
    For any monomial $m_{t}\in A^{\tilde{m}}_{{\hat{f}_{1}}}$, for every guess $r\in[D_{2}]$, we have $\lambda\equiv\mathcal{K}_{m_{t,r}}\mod p$, where $\mathcal{K}_{m_{t,r}}=D_{1}$. Thus, when considering the output of the automaton $\mathcal{M}$ on the monomial $m_{\ell}$ as the sum of two entries
@@
    Conversely, for the above fixed monomial $m_{\ell}\in B_{F_{1}}^{\tilde{m}}$, there exists a guess $r\in[D_{2}]$ such that $\mathcal{K}_{m_{\ell,r}}\neq D_{1}$ (by Claim 4.1.2). Therefore, this computation path will not reach either of the final states $q_{f_{1}}$ or $q_{f_{2}}$. Thus, the coefficient $\alpha_{m_{\ell}}$ of the monomial $m_{\ell}$ is scaled by at most a factor of $(D_{2}-1)$ in $M[q_{0},q_{f_{1}}]+M[q_{0},q_{f_{2}}]$.
@@
    $f_{\lambda}=D_{2}\cdot\hat{f}_{1}+\tilde{F}_{1},$